# Analyzing (In)Abilities of SAEs via Formal Languages

### v3

@@ 1 Introduction
In recent years, mechanistic interpretability has gained currency as an approach towards understanding the functioning of language models (LMs) Nanda et al. (2023); Olsson et al. (2022); Wang et al. (2022); Li et al. (2024a). A mechanistic interpretability paradigm that has seen remarkable progress especially is sparse autoencoders (SAEs), which aims to disentangle model representations into independent, interpretable components Lieberum et al. (2024); Bricken et al. (2023); Cunningham et al. (2023); Engels et al. (2024); Gao et al. (2024); Marks et al. (2024); Rajamanoharan et al. (2024) (see App. A for related work). Similar interpretability approaches based on autoencoders have been extensively studied in the vision domain (see Fig. 1), from both empirical and theoretical viewpoints Locatello et al. (2019); Träuble et al. (2021); Lachapelle et al. (2023); Locatello et al. (2020); Klindt et al. (2020). Specifically, using synthetic testbeds where a ground-truth set of latents underlying the data-generating process can be listed Higgins et al. (2017), these studies have established results on the (non-)identifiability of disentangled representations via autoencoders and argued for the importance of inductive biases, which motivate different regularization terms. We argue a similar study is needed for understanding the (in)abilities of SAEs for interpreting LM representations.
@@
  • We demonstrate, inline with results from vision, that identifiability of disentangled features is not robust (Sec. 3). Results are highly sensitive to changes in normalization methods, hyperparameters, and other such settings. For example, models trained under $L_{1}$ regularization consistently fail to find interpretable features; additionally, normalization has widely varying effects on the reconstruction capabilities of the SAEs.
@@ 2.2 SAE Architecture
where $W_{\text{enc}}\in\mathbb{R}^{h\times d}$, $W_{\text{dec}}\in\mathbb{R}^{d\times h}$, and $b_{\text{enc}}\in\mathbb{R}^{h},b_{\text{dec}}\in\mathbb{R}^{d}$. We pick $h$ from $\{d,2d,4d,8d\}$ and $\sigma=\text{ReLU}$ as the activation function. We refer to the encoder transform (up to and including the nonlinearity) as $E$, the decoder as $D$, and the hidden representation (the output of the encoder) as a latent. When these latents have interpretable explanations, identified correlationally (Sec. 3.1), we also refer to them as features.
@@ 3 Results
3.1 Do SAEs Find Features?
Qualitatively, we observe that top-$k$-regularized SAEs tend to have more interpretable features than $L_{1}$-regularized ones. The former results in interpretable features across all languages—we find features representing fundamental aspects of the corresponding grammars, as discussed next briefly. See App. C for further results, including other features we are able to identify and how strongly features correlate with their claimed explanations.
@@ 3.3 Does Correlation Imply Causation?
In the case of Expr, we have seen in Sec. 3.1 that there is a counter feature that activates when exactly one expression is left to complete the sequence. We pass an incomplete sequence to the model, and intervene on this feature by clamping it to one of $\{-v_{\text{max}},0,v_{\text{max}}\}$, where $v_{\text{max}}$ is the maximum absolute value of the latent (see Sec. 2.3). We then examine the generations that result in each case. We expect that high clamp values will cause the model to generate only one more expression (even if more may be required) and low values will cause it to generate more than one, or perhaps end the sequence immediately (even if exactly one is needed); or vice versa, since we cannot be certain of the direction to apply intervention in. Tab. 2 presents the results of this experiment. For each input sample, we mention the number of expressions required to complete it, and the number of expressions actually generated by the model under each intervention. We do not see the expected—or indeed any—causal effect of intervening on this feature—it appears that this feature is only correlational in nature.
Similarly, we examine our English grammar. Here, we examine the effects of intervening on the part-of-speech features described in Sec. 3.1. Note that for a feature correlating with, say, adjectives, we can expect that if it has a causal effect, it must control the next-token distribution for that part-of-speech (as this is the only task the model is trained on). We thus apply our causality protocol (Sec. 2.3) to the feature correlating with adjectives. The results of this experiment are shown in Fig. 5. In this case, as in Expr, we do not see a causal effect of intervening on this feature. See App. D for similar results on other parts-of-speech.
@@ 4 Incentivizing Causality
Our results above indicate that though SAEs are competent at identifying features correlated with semantically meaningful concepts, these features may not have the expected (or any) causal effects. This mirrors similar observations made in vision, where it has been noted (Locatello et al., 2019) that mere data reconstruction pipelines (e.g., based on autoencoders) can fail to disentangle the latent factors of a generative model. However, later work on the topic demonstrated that correlations in the data-generating process (e.g., correlation in nearby frames of a video) can be leveraged to obtain disentangled representations (Locatello et al., 2020; Klindt et al., 2020). As this method relies on supervision via other data samples, rather than a ground truth, it is described as weak supervision. A warranted question then is whether weak supervision can be elicited in language modeling scenarios too.
@@ 4.2 Nature of Causal Features
As in the SAEs trained without this loss (Sec. 3.1), we identify a number of features correlating with parts of speech—specifically, adjectives, verbs and adverbs. Now, we predict that if these features have a monotonic causal function, they shift the output logit distribution towards the one predicted by the corresponding part of speech. For example, a feature correlating with adjectives should promote the probability of nouns being predicted next, as nouns are the only PoS allowed to come after adjectives (see App. B for the exact grammar). In the case of verbs, anything may appear as the next token except another verb; thus we expect the probability of verbs to be downweighted by this feature.
Fig. 7 presents the effect of these interventions on the PoS distribution across sentences, along with the grammaticality of generations. The effect is as we hypothesized. We also note that as soon as the distribution significantly shifts away from the uncorrupted distribution, the fraction of grammatical generations drops drastically. This supports, as pointed out in Träuble et al. (2021); Ahmed et al. (2020), the difficulty of learning disentangled representations from correlated data (in this case, with respect to PoS distribution and grammaticality). More examples of the causal function of these features can be found in App. D.3.
@@ 4.3 Inductive Biases in our Pipeline
Overall, inline with prior work on autoencoders-based interpretability (Locatello et al., 2019), our results highlight the importance of recognizing the inductive biases of any SAE paradigm for interpretability. We expect that any method aiming to integrate causality into feature identification will similarly have inductive biases, and no single method will overcome all the limitations of the paradigm.
@@ 6 Limitations
ESL conceived the project direction and defined the broader project narrative around disentangled representation learning. ESL wrote code for the data generation process and model training / evaluation. AM led the feature interpretability analysis, with several inputs from ESL. ESL and AM co-wrote the paper. ESL devised the causal loss idea for Section 4, and AM implemented the method and qualitatively analyzed the results. DSK acted as primary senior advisor, with inputs from MS.
@@ A.5 Synthetic Data in Interpretability
Formal languages have also supplied synthetic data, particularly to understand the expressivity of various sequence architectures. We have discussed in the main paper (Sec. 3.1) the results of Bhattamishra et al. (2020), who use the languages of Dyck-1, Shuffle-Dyck and boolean expressions to establish the role of attention in generalization across sequence lengths, and contextualize the set of languages recognizable by transformers within the Chomsky hierarchy (Sipser, 1996).
@@ Appendix B Formal Grammars
We use, as mentioned in Sec. 2, three formal grammars on which we train transformer-based language models. The exact specification of these grammars, in terms of a context-free grammar (CFG), is presented below. Note that the actual data generation process is defined by a probabilistic CFG, which requires probabilities to be assigned to each of the productions of a nonterminal. We omit these probabilities for legibility here, but readers can refer to our codebase (https://anonymous.4open.science/r/pcfg-sae-mint24-4A64) for details.
@@ B.3 English
Each part of speech – nouns (N), verbs (V), pronouns (Pro), conjunctions (Conj), adjectives (Adj) and adverbs (Adv) – have ten tokens each. We omit the production rules listing these for brevity, but refer readers to the codebase (https://anonymous.4open.science/r/pcfg-sae-mint24-4A64) for details.
@@ Appendix C Features
We describe here several other features we were able to identify in our SAEs. Results are shown in Tab. 3. Note that all the SAEs listed here are trained without pre_bias and without normalization. Thus, the hyperparameters specified in column 3 are the expansion factor (ratio between hidden size and input size) and, according to the regularization method, either the $L_{1}$ coefficient or $k$. For each feature, we note the hyperparameter settings of the SAE, our hypothesized explanation, and the correlation (Pearson coefficient) between the feature’s activation and the explanation.
@@ Appendix D Causality of Features
We present here a detailed outline of our experimental protocols in the causality experiments. We also examine other features for causal effects, particularly the counter feature for Expr (described in Sec. 3), and part-of-speech features for English. In the latter case, we examine features identified by ordinary SAEs and SAEs with causal loss in Sec. D.2 and Sec.D.3 respectively. We refer the reader to our codebase for causal experiments (https://anonymous.4open.science/r/pcfg-sae-causal-arr-oct24-C607) for the implementation of these protocols.
@@ D.2 English: Parts of Speech (without causal loss)
The results of interventions on various features are shown in Fig. 9. For each part of speech that we show, we follow the same process as in the case of adjectives (Fig. 5):
@@ Appendix E Evidence for Linearity Incentive
Figure 9: Behavior of the English model under interventions. We examine the causality of features corresponding to other parts of speech: verbs ((a) and (b)), conjunctions ((c) and (d)), and adverbs ((e) and (f)). At each clamp value (x-axis), we plot the distribution of nouns, pronouns, adjectives, verbs, adverbs, and conjunctions.
@@ Appendix G Training Details
Figs. 12 to 19 show the details of the various hyperparameter settings we have considered for SAEs, along with plots of the losses during training. For convenience, we record these details in Tab. 5 as well. Note that, in the case of $L_{1}$-regularized SAEs, the training loss is the sum of the reconstruction loss (the MSE loss between the SAE reconstructions and the inputs) and the regularization loss (the norm of the hidden representation, multiplied by the $L_{1}$ coefficient). The SAEs are trained for 5000 iterations, with losses logged every 25 iterations.

### v4

@@ 1 Introduction
In recent years, mechanistic interpretability has gained currency as an approach towards understanding the functioning of language models (LMs) Nanda et al. (2023); Olsson et al. (2022); Wang et al. (2022); Li et al. (2024a). A mechanistic interpretability paradigm that has seen remarkable progress especially is sparse autoencoders (SAEs), which aims to disentangle model representations into independent, interpretable components Lieberum et al. (2024); Bricken et al. (2023); Cunningham et al. (2023); Engels et al. (2024); Gao et al. (2024); Marks et al. (2024); Rajamanoharan et al. (2024) (see App. A for related work). Similar interpretability approaches based on autoencoders have been extensively studied in the vision domain (see Fig. 1), from both empirical and theoretical viewpoints Locatello et al. (2019); Träuble et al. (2021); Lachapelle et al. (2023); Locatello et al. (2020); Klindt et al. (2020). Specifically, using synthetic testbeds where a ground-truth set of latents underlying the data-generating process can be listed Higgins et al. (2017), these studies have established results on the (non-)identifiability of disentangled representations via autoencoders and argued for the importance of inductive biases, which motivate different regularization terms. For instance, Locatello et al. (2019) prove that for any given dataset (modeled as a multivariate distribution), there are infinitely many generative models, all entangled with each other, and remark that inductive biases are necessary to “select” from among these distributions. Träuble et al. (2021) show that in a correlated dataset (which real-world datasets tend to be), a perfectly disentangled model would lead to suboptimal log-likelihood, and therefore would not be learned without significant inductive biases. We argue a similar study is needed for understanding the (in)abilities of SAEs for interpreting LM representations.
@@
  • We demonstrate, in line with results from vision, that identifiability of disentangled features is not robust (Sec. 3). Results are highly sensitive to changes in normalization methods, hyperparameters, and other such settings. For example, models trained under $L_{1}$ regularization consistently fail to find interpretable features; additionally, normalization has widely varying effects on the reconstruction abilities.
@@ 2.2 SAE Architecture
where $W_{\text{enc}}\in\mathbb{R}^{h\times d}$, $W_{\text{dec}}\in\mathbb{R}^{d\times h}$, and $b_{\text{enc}}\in\mathbb{R}^{h},b_{\text{dec}}\in\mathbb{R}^{d}$. We pick $h$ from $\{d,2d,4d,8d\}$ and $\sigma=\text{ReLU}$ as the activation function. We refer to the encoder transform (up to and including the nonlinearity) as $E$, the decoder as $D$, and the hidden representation (the output of the encoder) as a latent. When these latents have interpretable explanations, identified correlationally (Sec. LABEL:sec:features), we also refer to them as features.
@@ 3 Results
Qualitatively, we observe that top-$k$-regularized SAEs tend to have more interpretable features than $L_{1}$-regularized ones. The former results in interpretable features across all languages—we find features representing fundamental aspects of the corresponding grammars, as discussed next briefly. See App. C for a description of our feature identification pipeline, and further results, including other features we are able to identify and how strongly features correlate with their claimed explanations.
@@ 3.2 Does Correlation Imply Causation?
In the case of Expr, we have seen in Sec. LABEL:sec:features that there is a counter feature that activates when exactly one expression is left to complete the sequence. We pass an incomplete sequence to the model, and intervene on this feature by clamping it to one of $\{-v_{\text{max}},0,v_{\text{max}}\}$, where $v_{\text{max}}$ is the maximum absolute value of the latent (see Sec. 2.3). We then examine the generations that result in each case. We expect that high clamp values will cause the model to generate only one more expression (even if more may be required) and low values will cause it to generate more than one, or perhaps end the sequence immediately (even if exactly one is needed); or vice versa, since we cannot be certain of the direction to apply intervention in. Tab. 2 presents the results of this experiment. For each input sample, we mention the number of expressions required to complete it, and the number of expressions actually generated by the model under each intervention. We do not see the expected—or indeed any—causal effect of intervening on this feature—it appears that this feature is only correlational in nature.
Similarly, we examine our English grammar. Here, we examine the effects of intervening on the part-of-speech features described in Sec. LABEL:sec:features. Note that for a feature correlating with, say, adjectives, we can expect that if it has a causal effect, it must control the next-token distribution for that part-of-speech (as this is the only task the model is trained on). We thus apply our causality protocol (Sec. 2.3) to the feature correlating with adjectives. The results of this experiment are shown in Fig. 5. In this case, as in Expr, we do not see a causal effect of intervening on this feature. See App. D for similar results on other parts-of-speech.
@@ 4 Incentivizing Causality
Our results above indicate that though SAEs are competent at identifying features correlated with semantically meaningful concepts, these features may not have the expected (or any) causal effects. This mirrors similar observations made in vision, where it has been noted that mere data reconstruction pipelines (e.g., based on autoencoders) can fail to disentangle the latent factors of a generative model (Locatello et al., 2019). However, later work on the topic demonstrated that correlations in the data-generating process (e.g., correlation in nearby frames of a video) can be leveraged to obtain disentangled representations (Locatello et al., 2020; Klindt et al., 2020). As this method relies on supervision via other data samples, rather than a ground truth, it is described as weak supervision. A warranted question then is whether weak supervision can be elicited in language modeling scenarios too.
@@ 4.2 Nature of Causal Features
As in the SAEs trained without this loss (Sec. LABEL:sec:features), we identify a number of features correlating with parts of speech—specifically, adjectives, verbs and adverbs. Now, we predict that if these features have a monotonic causal function, they shift the output logit distribution towards the one predicted by the corresponding part of speech. For example, a feature correlating with adjectives should promote the probability of nouns being predicted next, as nouns are the only PoS allowed to come after adjectives (see App. B for the exact grammar). In the case of verbs, anything may appear as the next token except another verb; thus we expect the probability of verbs to be downweighted by this feature.
Fig. 7 presents the effect of these interventions on the PoS distribution across sentences, along with the grammaticality of generations. The effect is as we hypothesized. We also note that as soon as the distribution significantly shifts away from the uncorrupted distribution, the fraction of grammatical generations drops drastically. This supports, as pointed out in Träuble et al. (2021); Ahmed et al. (2020), the difficulty of learning disentangled representations from correlated data (in this case, with respect to PoS distribution and grammaticality). In fact, recent work Bhalla et al. (2024); Wu et al. (2025) observes that intervention-based methods (in comparison with probing-based methods) suffer from the drawback of damaging models’ output quality in terms of coherence, creating a control-capabilities tradeoff. This indicates that the problem is possibly more general than simple semantic disentanglement, and this is a valuable avenue for future research. We note, furthermore, that this is another validation of our setup (like the scaling law described in Sec. LABEL:sec:features), as it demonstrates behaviors also observed in more complex, natural language settings. More examples of the causal function of these features can be found in App. D.3.
@@ 4.3 Inductive Biases in our Pipeline
Overall, in line with prior work on autoencoders-based interpretability (Locatello et al., 2019), our results highlight the importance of recognizing the inductive biases of any SAE paradigm for interpretability. We expect that any method aiming to integrate causality into feature identification will similarly have inductive biases, and no single method will overcome all limitations of the paradigm.
@@ 6 Limitations
ESL conceived the project direction and defined the broader project narrative around disentangled representation learning. ESL wrote code for the data generation process and model training / evaluation, with inputs from AM in expanding to several languages. AM led the feature interpretability analysis, with several inputs from ESL. ESL and AM co-wrote the paper. ESL devised the causal loss idea for Section 4, and AM implemented the method and qualitatively analyzed the results. DSK acted as primary senior advisor, with inputs from MS.
@@ A.5 Synthetic Data in Interpretability
Formal languages have also supplied synthetic data, particularly to understand the expressivity of various sequence architectures. We have discussed in the main paper (Sec. LABEL:sec:features) the results of Bhattamishra et al. (2020), who use the languages of Dyck-1, Shuffle-Dyck and boolean expressions to establish the role of attention in generalization across sequence lengths, and contextualize the set of languages recognizable by transformers within the Chomsky hierarchy (Sipser, 1996).
@@ Appendix B Formal Grammars
We use, as mentioned in Sec. 2, three formal grammars on which we train transformer-based language models. The exact specification of these grammars, in terms of a context-free grammar (CFG), is presented below. Note that the actual data generation process is defined by a probabilistic CFG, which requires probabilities to be assigned to each of the productions of a nonterminal. We omit these probabilities for legibility here, but readers can refer to our (https://github.com/Abhinav271828/pcfg-sae-causal-arr-oct24) for details.
@@ B.3 English
Each part of speech – nouns (N), verbs (V), pronouns (Pro), conjunctions (Conj), adjectives (Adj) and adverbs (Adv) – have ten tokens each. We omit the production rules listing these for brevity, but refer readers to the (https://github.com/Abhinav271828/pcfg-sae-causal-arr-oct24) for details.
@@
Empty stack of type 0. Dyck (8, $k=128$) 0.925
Empty stack of type 1. Dyck (8, $k=128$) 0.814
@@ Appendix C Features
We describe here the methodology of finding features in SAEs, and list several other features we were able to identify.
C.1 Feature Identification Pipeline
C.2 Experimental Protocols: Feature Identification
We avoid the automated interpretability paradigm outlined in Bills et al. (2023), in order to avoid concerns of LLMs being unfamiliar with the specific formal grammars we use. Instead, in line with Bricken et al. (2023), we begin with manual inspection of each feature, and hypothesize potential explanations for its semantics using positive and negative samples. As mentioned in the same work, this must also be informed by the reconstruction capabilities of the SAE—we therefore restrict our analysis to SAEs that are able to achieve high reconstruction accuracy (defined in 3.1. Following Makelov et al. (2024), we compare the activations of the latents with the predictions of the explanations. We use the correlation between the two as a score for the quality of the explanation. Marks et al. (2024) use a manual interpretability pipeline similar to ours, with the added information of the latent’s causal role (through ablations). However, our aim is primarily to evaluate the quality of purely correlational explanations, which many feature identification pipelines (both manual and automated) rely upon. The causal nature of the explanations is explored in Sec. 3.2
The manual inspection is based on an a priori notion of the latent generative factors inherent in each formal language, which we briefly outline. In Dyck-2, we expect some computation to be based on the depth of the current token (i.e., the number of unclosed brackets preceding it). In Expr, we expect a counter that keeps track of the number of expressions yet to be generated. In English, we expect features corresponding to each part of speech in the grammar. These are all variables both necessary and sufficient for the data-generation process. We note, however, that these are not the only possibilities—each language may have several equivalent reformulations, but we believe these are the most intuitive and suitable for the autoregressive decoding paradigm. We discuss in more detail the exact formulation of the grammars and their relation to these generative factors in App. B.
C.3 Identified Features
Results are shown in Tab. 3. Note that all the SAEs listed here are trained without pre_bias and without normalization. Thus, the hyperparameters specified in column 3 are the expansion factor (ratio between hidden size and input size) and, according to the regularization method, either the $L_{1}$ coefficient or $k$. For each feature, we note the hyperparameter settings of the SAE, our hypothesized explanation, and the correlation (Pearson coefficient) between the feature’s activation and the explanation.
@@ Appendix D Causality of Features
We present here a detailed outline of our experimental protocols in the feature identification and the causality experiments. We also examine other features for causal effects, particularly the counter feature for Expr (described in Sec. 3), and part-of-speech features for English. In the latter case, we examine features identified by ordinary SAEs and SAEs with causal loss in Sec. D.2 and Sec.D.3 respectively. We refer the reader to our for causal experiments (https://github.com/Abhinav271828/pcfg-sae-causal-arr-oct24) for the implementation of these protocols.
@@ D.2 English: Parts of Speech (without causal loss)
The results of interventions on various features are shown in Fig. LABEL:fig:interventions_app. For each part of speech that we show, we follow the same process as in the case of adjectives (Fig. 5):
@@ Appendix G Training Details
Figs. 11 to 18 show the details of the various hyperparameter settings we have considered for SAEs, along with plots of the losses during training. For convenience, we record these details in Tab. 5 as well. Note that, in the case of $L_{1}$-regularized SAEs, the training loss is the sum of the reconstruction loss (the MSE loss between the SAE reconstructions and the inputs) and the regularization loss (the norm of the hidden representation, multiplied by the $L_{1}$ coefficient). The SAEs are trained for 5000 iterations, with losses logged every 25 iterations.